OpenStudy (anonymous):

What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5 these are the answer chioces a.(−10, −5) b.(−7, −2) c.(7, 2) d.(5, 10)

OpenStudy (anonymous):

Do you want X and Y? Those choices makes no sense if it is X and Y you want to find. I get X = 43/4 and Y = 64/4...